OpenStudy (anonymous):

Find the sum of the infinite geometric series. 1 + 1/6 + 1/36+ 1/216 +...

OpenStudy (anonymous):

\[S = \frac{a}{1- r}\] a = 1 r = 1/6 S = 1 / (1 - 1/6) S = 1 / (5/6) S = 6/5

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Wait how did you find r?